40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 12 00:40:23 2014, Jenne, Update, LSC, DRMI locking 
    Reply  Fri Sep 12 11:25:42 2014, Koji, Update, LSC, DRMI locking 
       Reply  Fri Sep 12 12:00:59 2014, ericq, Update, LSC, DRMI locking 
    Reply  Fri Sep 12 14:11:17 2014, ericq, Update, LSC, DRMI locking 7x
       Reply  Wed Sep 17 15:40:00 2014, ericq, Update, LSC, DRMI locking dcAngleDiff_srcL.pdfdrfpmiVertexSensing.zip
          Reply  Tue Sep 23 17:37:10 2014, ericq, Update, LSC, DRMI locking SRCcorrection.pdf
             Reply  Wed Sep 24 11:02:38 2014, manasa, Update, LSC, Moving SRM 
                Reply  Thu Sep 25 11:33:41 2014, Jenne, Update, LSC, POY alignment laser 
Message ID: 10538     Entry time: Thu Sep 25 11:33:41 2014     In reply to: 10531
Author: Jenne 
Type: Update 
Category: LSC 
Subject: POY alignment laser 

Quote:

I looked at the CAD layout and it seems like we will clearly be clipping POY if we move SRM by 7.5cm. Since POY is not visible at low power, we cannot be sure about the clipping.

 I was bad and forgot to elog this yesterday (bad grad student!), but I setup a laser pointer to show us where the POY beam is. 

To do this, I removed the tiny mirror that sends the beam to the POY RF PD (so we do not have POY to lock the Yarm right now.  I think Q has successfully been using AS).  The laser pointer goes through 2 temporary steering mirrors, then passes through the place that the tiny mirror usually sits, and then travels along the POY path into the vacuum system.  The idea here is that we should be able to adjust the laser pointer and the temp steering mirrors, and not touch any of the actual POY mirrors, but still get the green beam to go all the way to ITMY.  Yesterday I confirmed that the laser pointer was hitting the in-vac POY pickoff mirror, and today Q and Manasa are doing final adjustment to get the beam all the way to the ITM. 

ELOG V3.1.3-